Two planes are 60 kilometers apart. Flying towards each other, they meet in 10 minutes. When flying in the same direction (at the same speeds as before), the faster plane overtakes the slower plane in 30 minutes. What is the speed of each plane in kilometers per hour?
Answer by mananth(16949)

Plane A speed be x
Plane B speed = y
Towards each other -combined speed = x+y
distance = 60 km
Time = 10/60 = 1/6 hours
t=d/r
1/6 = 60/(x+y)
cross multiply
x+y=360-------------1
same direction
catch up speed = x-y
catch up time = 1/2 hour
D= 60 km
1/2= 60/(x-y)
x-y=120.............2
Add equations (1) & (2)
2x=480
x==240 km/h speed of A
360-240=120 km/h speed of B